if the sum of the measures of the interior angles of a polygon is 1980 degrees, how many sides does it have?
How do I do this? Is there a formula?
Thanks!

Answer by macston(5194)

.
n=number of sides; S=sum of interior angles
.
S = 180(n-2)
S/180=n-2
(S/180)+2=n
.
(1980/180)+2=n
11+2=n
13=n
.
ANSWER: The polygon has 13 sides.
.